OpenStudy (dragontears1128):

A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a/2=ae^−3.392t

OpenStudy (slenderman):

hold up im looking for a teacher on this one

OpenStudy (slenderman):

e^(-6.345t)=1/2 t=0.11 days ok so my teacher wrote this down on my paper, it dosent make sense to me @AceSpeedFighter do you know this?

OpenStudy (slenderman):

@TheSmartOne

OpenStudy (slenderman):

@devincoolage1001 do you know this

OpenStudy (sweetburger):

Are we assuming the variable a is not representing the initial value on the right hand side of the equation and that a is not representing the final value in mass of the substance on the left hand side?

OpenStudy (sweetburger):

If a is the same variable on both sides of the equation then you could actually solve for t. However, that doesnt seem likely given the question appears to be missing information.

OpenStudy (dragontears1128):

A radio active isotope decays according to the exponential decay equation where t is in days. For the half life: The half life is the solution (t) of the equation : a/2=ae^−3.392t Round to the thousandths place

OpenStudy (dragontears1128):

@sweetburger